A candy store sells boxes of candy containing caramels and chocolates. Each box sells for $12.50 and holds 30 pieces of candy (all pieces are the same size). If the caramels cost $0.25 to produce and the chocolates cost $0.45 to produce, how many of each should be in a box to make a profit of $3.00?
Answer by ptaylor(2198)


Let x= number of caramels in the box
And 30-x=number of chocolates in the box
Now we are told that the cost of the caramels ($0.25x) plus the cost of the chocolates ($0.45(30-x)) must equal $12.50-$3.00=$9.50 in order to give us a profit of $3.00. So that gives us:

$0.25x+$0.45(30-x)=$9.50 multiplying out
$0.25x+$13.5-$0.45x=$9.50 subtract $13.50 from both sides
$0.25x+$13.50-$13.50-$0.45x=$9.50-$13.50 collecting like terms
-$0.20x=-$4
x=20-------------number of caramels in the box
30-x=30-20=10--------------number of chocolates in the box
CK
$0.25(20)+$0.45(10)=$12.50-$3.00
$5+$4.50=$9.50
$9.50=$9.50
Hope this helps-----ptaylor